OpenStudy (anonymous):

3.05, 4.08]A company builds Electronic Entertainment Centers. It costs $8,700 to build 10 Electronic Entertainment Centers and $14,300 to build 20 Electronic Entertainment Centers. Which equation models the cost, C(x), as a linear function of the number of Electronic Entertainment Centers, x? C(x) = 3100x − 560 C(x) = 560x − 3100 C(x) = 3100x + 560 C(x) = 560x + 3100 Help Please!! ~Dylan

OpenStudy (anonymous):

c(x)= 560x+ 3100 put values of x number of units product you will have answer

OpenStudy (anonymous):

How did you get that equation?

OpenStudy (anonymous):

put value od x c= 560*10 +3100 = 8700 c=560*20+3100 = 14700

OpenStudy (anonymous):

Ohhh! ok thanks sooo much :)

OpenStudy (anonymous):

:)